A manufacturer of power tools claims that the average amount of time required to assemble their top-of-the-line table saw is 80 minutes with a standard deviation of 40 minutes. Suppose a random sample of 64 purchasers of this table saw is taken. The probability that the sample mean will be greater than 88 minutes is __________. What type of problem is this?

Poisson

Random variable

Sample mean

Sample proportion

Normal
Answer #1

MEAN= 80 MINUTES AND S.D= 40 MINUTES

n=64

sample mean will be 80 minutes and S.D(= 40/sqrt(64)= 40/8=5

P ( X>88 )=P ( X−μ>88−80 )=P ((X−μ)/σ>(88−80)/5)

Since Z=(x−μ)/σ and (88−80)/5=1.6 we have:

P ( X>88 )=P ( Z>1.6 )

Use the standard normal table to conclude that:

P (Z>1.6)=0.0548

This problem is NORMAL. But actually it is sampling distribution problem.